# BihoT: A Large-Scale Dataset and Benchmark for Hyperspectral Camouflaged Object Tracking

### v1

@@ V-C Comparison with State-of-the-arts
To demonstrate the ability to handle different challenge scenarios of our proposed SPDAN, we conduct attribute-based experiments with the selected SOTA trackers on the HOTC-2020 and Bihot datasets. Fig. 10 shows the results of the HOTC-2020 dataset. For the success plot, SPDAN achieves the best results in 8 out of all 11 scenarios. This proves the effectiveness of our method to handle the complex scenarios. The SPDAN performs better in scenarios where visual features are damaged, including the loss of color, which is particularly evident in the IV and LR scenarios. In these cases, the color representation of the object is compromised, leading to a significant reduction in the availability of the visual information. Despite these challenges, our proposed SPDAN yields an AUC of 0.703, which is 9.4% higher than the second-ranked SENSE in the IV scenario, indicating the importance of spectral information in extracting robust features. Additionally, the deformation of the object presents another critical challenge in feature extraction. In scenarios such as DEF and SV, changes in the object’s shape or appearance can severely impact the ability to track it accurately, as standard visual features become unreliable. Our method achieves the best results in both metrics, proving that it can effectively extract cross-modality spectral features and has strong robustness to object deformation. Our method also demonstrates an excellent ability to deal with occluded scenes. SPDAN achieves 0.708 of the AUC score and 0.954 of the DP_20, which is 5.8% and 3.5% higher than the second-ranked method. This proves that our proposed DAM can effectively handle frequent short-term occlusion.
In addition, we also conduct attribute-based experiments on the proposed BihoT dataset. The experimental results are listed in Table IX. It is worth noting that to make more appropriate attribute comparisons, we redefine the challenge scenarios based on the characteristics of the BihoT, as detailed in Table II. Unlike the settings in HOTC-2020, we remove attributes (i.e., DEF, MB, and OV) and introduce scenarios such as SC (Spectral Consistency) and SV (Spectral Variability) to more accurately assess the utilization of spectral information. Consequently, the number of attributes evaluated is reduced from 11 to 9. It can be seen that our SPDAN exhibits superior performance in different scenarios, which further proves the effectiveness of our proposed method. Specifically, in the two spectral-related scenarios SC and SV, SPDAN achieves AUC of 0.627 and 0.702 and DP of 0.789 and 0.830. The proposed method can accurately predict the position of the object in the presence of visual similarity, spectral differences, and spectral similarity interference, demonstrating that our spectral embedding network and prompt-based learning method can accurately extract spectral features for the accurate tracking.
@@ VI Conclusion
In this paper, we proposed a hyperspectral camouflaged object tracking task and constructed a large-scale HCOT dataset, BihoT, which contains 41,912 hyperspectral images where objects have similar appearances, diverse spectrums, and frequent occlusion. Then, we proposed a spectral prompt-based distractor-aware network (SPDAN) as a simple yet effective baseline, which combines discriminative information from the HS and RGB modalities through prompt-based learning. A spectral embedding network was designed to transform the spatial and spectral information into tokenized features. To make the tracking results more robust, we introduced a distractor-awareness module to perceive the distractor and rectify the original results. The quantitive and qualitative experimental results demonstrated the effectiveness of our method.

### v2

@@ V-C Comparison with State-of-the-arts
To demonstrate the ability to handle different challenge scenarios of our proposed SPDAN, we conduct attribute-based experiments with the selected SOTA trackers on the BihoT datasets. The experimental results are listed in Table XIII. It is worth noting that to make more appropriate attribute comparisons, we redefine the challenge scenarios based on the characteristics of the BihoT, as detailed in Table II. Unlike the settings in HOTC-2020, we remove attributes (i.e., DEF, MB, and OV) and introduce scenarios such as SC (Spectral Consistency) and SV (Spectral Variability) to more accurately assess the utilization of spectral information. Consequently, the number of attributes evaluated is reduced from 11 to 9. It can be seen that our SPDAN exhibits superior performance in different scenarios, which further proves the effectiveness of our proposed method. Specifically, in the two spectral-related scenarios SC and SV, SPDAN achieves an AUC of 0.627 and 0.702, and DP_20 of 0.789 and 0.830. The proposed method can accurately predict the position of the object in the presence of visual similarity, spectral differences, and spectral similarity interference, demonstrating that our spectral embedding network and prompt-based learning method can accurately extract spectral features for accurate tracking.
@@ VI Conclusion
In this paper, we proposed a hyperspectral camouflaged object tracking task and constructed a large-scale HCOT dataset, BihoT, which contains 41,912 hyperspectral images where objects have similar appearances, diverse spectrums, and frequent occlusion. Then, we proposed a spectral prompt-based distractor-aware network (SPDAN) as a simple yet effective baseline, which combines discriminative information from the HS and RGB modalities through prompt-based learning. A spectral embedding network was designed to transform the spatial and spectral information into tokenized features. To make the tracking results more robust, we introduced a distractor-awareness module to perceive the distractor and rectify the original results. The quantitive and qualitative experimental results demonstrated the effectiveness of our proposed method. In future work, more attention will be given to leveraging temporal information in video sequences to enhance HOT performance. By integrating time-series data, trackers may better capture object motion patterns and improve robustness in challenging scenarios such as occlusions or rapid appearance changes.